OpenStudy (anonymous):

m -- 16-9=-8 how to solve tihs??!

OpenStudy (anonymous):

it is a -8

OpenStudy (alfie):

sorry, misread. Is -8 so it's -56! :) m = -56

OpenStudy (anonymous):

thank you so u multiplied both sides by 8

OpenStudy (alfie):

I multiplied both sides by 7. \[\huge \frac{m}{7}*7 = - 8*7 \to m = - 56\]